Minutes of the 12th SOHO SWT Meeting

Portsmouth, England 15–16 June 1994 Contents

1 Action Items and Resolutions SWT 12 3 2 Agree Agenda and Actions Revision 3 3 Science Operations Report 4 4 EOF 4 5 Status of the Project 4 6 Status of Experiments 5 7 SOHO Intercalibration Committee report 5 8 SPWG report 5 9 Data Rights and Obligations 5 10 SOHO Data Archive 6 11 European SOHO Data Archive 6 12 Early Operations 6 13 Jitter Timeline 6 14 Guest Investigators Programme 6 15 SOP 7 16 Workshops and Conferences 7 17 Science Working Team executive meeting 7 Annex 1: Agenda 8 Annex 2: List of Participants SWT-12 9 Annex 3: Science Operations Report (V. Domingo) 11 Annex 4: EOF Layout (A. Poland) 14 Annex 5: SOHO CD-ROMs Distribution List (A. Poland) 15 Annex 6: Project Status (F. Felici) 20 Annex 7: SOHO Engineering Review (F. Vandenbussche) 27 Annex 8: Payload & Operations Development Status (C. Berner) 38 Annex 9: Cleanliness Review (R. Thomas) 74 Annex 10: Task Force Report (L. Culhane) 90 Annex 11: Status of Experiments 102

SOI-MDI (P. Scherrer) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 GOLF (A. Gabriel) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 VIRGO (C. Fröhlich) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 LASCO (G. Brueckner) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128 CDS (R. Harrison) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 SUMER (K. Wilhelm) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 SWAN (J.-L. Bertaux) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 CELIAS (D. Hovestadt) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 COSTEP (R. Müller-Mellin) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 ERNE (J. Torsti) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 EIT (J.-P. Delaboudiniere) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 UVCS (J. Kohl) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 216

Annex 12: SIC Report (J. Kohl) 232

Annex 13: SPWG Report (A. Gabriel) 235

Annex 14: Data Rights and Obligations (V. Domingo) 240

Annex 15: SOHO Data Archive (L. Sanchez) 244

Annex 16: European SOHO Data Archive (V. Domingo) 254

Annex 17: Early Operations (P. Martens) 261

Annex 18: IACG Workshop Report (V. Domingo) 269

1 Action Items and Resolutions SWT 12

Action 12.1 on PI’s: To comment and update the early operations timeline plan (TTP + commissioning phase [first 30 days after HOI] + first 30 days of nominal operations). Input to P. Martens before July 31, 1994

Action 12.2 on Project: To provide the S/C commissioning plan before August 31, 1994.

Action 12.3 on PI’s: To define and provide their team’s data rights and publication policy (to be included in the SOP as Annex E) and comment on Table 3.7 of SOP Draft issue 2.0 (May 1994) before July 31, 1994.

Action 12.4 on PI’s: To submit camera-ready manuscript of instrument paper for the SOHO special issue in SOLAR PHYSICS. Input to B. Fleck before October 31, 1994.

Resolution on European SOHO Data Archives

The SOHO SWT endorses the ESA plan to seek institutions interested in hosting the SOHO data in Europe. The European archive(s) will be a copy of the EOF archive (data library) and adhere to the same access rules established for the EOF archive.

2 Agree Agenda and Actions Revision

Agenda — annex 1 List of Participants — annex 2

Actions revision

Action 11.1 on PI’s: Status: closed To provide plans/wishes for cruise phase activities (first 145 days). Input to PS before Feb. 14, 1994.

Action 11.2 on PI’s: Status: closed To fill out their part of Table I in data rights document. Deadline: Feb. 14, 1994. (This Table will be included in the next SOP. Without PI input it will be assumed the PS is trusted to fill out the table reasonably.)

Action 11.3 on PI’s: Status: closed To comment on the first draft of the SOHO GI programme AO, before Feb. 14, 1994. In particular we need the PI Team’s view on how they see GI participation for their instrument, since the “30% of observing time” criterion is meaningless or even unreasonable for some instruments.

Action 11.4 on PI’s: Status: open for GOLF, VIRGO, SOI, EIT,

LASCO, SWAN, COSTEP, ERNE

To write up and distribute the “Instrument Science Report” (e.g. CDS “Blue Book”) before July 1, 1994.

Action 11.5 on PI’s: Status: closed To provide input for Annex A (list of “Institutions involved in data processing and analysis”) of SOP before Feb. 14, 1994.

3 Science Operations Report

V. Domingo presented the schedule for the EOF/EAF commissioning and science operations (see annex 3). G. Brueckner expressed his concerns about NASA funding for MODA. He mentioned that with his present resources he cannot hire a single person to move to the EOF and therefore cannot meet this “ambitious” schedule.

V. Domingo announced that members of the Project Scientist Office (PSO) will visit the PI teams to discuss and coordinate the planning and commanding software.

4 EOF

A. Poland presented the layout of the EOF (annex 4) and a preliminary distribution list of SOHO CD-ROMs (annex 5). World Wide Web forms have been selected as the front-end to the SOHO online catalogs because it is easy-to-use, uses a client-server approach, and the client software is readily available (Mosaic). For this reason as well as for quick and easy access to SOHO information on the SOHO Web pages it is strongly recommended that everybody familiarizes with Mosaic.

SOHO information on Internet: Mosaic: URL: http://umbra.gsfc.nasa.gov Anonymous ftp: ftp umbra.gsfc.nasa.gov

login: anonymous

passwd: <e-mail address> cd /pub/soho

5 Status of the Project

a) Overall FM PLM AIV schedule (F. Felici) see handout — annex 6

b) SOHO engineering review (F.C. Vandenbussche) see handout — annex 7

c) Payload & operations development status (C. Berner) see handout — annex 8

– PLM delivery do MMS-F by mid of August

– System AIV test beginning of Sep. 94

– SOHO AIV programme to be completed 3 April 95

– The 30 days commissioning phase after HOI will be clarified clarify by Project and scientists (see action 12.2)

– Project emphasized that to meet the programme deadline experimenters shall:

• deliver flightworthy experiments in agreed time slots

• support AIV system activities as required by planned test activities

• provide thermal analytical models and TB/TV inputs according to the timeline defined d) Cleanliness (R. Thomas) see handout — annex 9 D. Michels expressed serious concerns that the time allocated to TB/TV test is not adequate, due to schedule pressure. He requested criteria for outgassing, pressure etc. before TB/TV tests be performed.

6 Status of Experiments

In a special session held on Wednesday, June 15, the PI’s presented the status of their experiments to DSci, R. Bonnet (for hardcopies of the PI presentations see annex 11). The session was opened by L. Culhane who gave a summary of the Task Force activities (cf. annex 10). After the PI presentations there was a visit to Matra Marconi where one could see the FM PLM. DSci expressed serious concerns on the viability of the SOHO mission if contamination of instruments was not better controlled by the PI’s. In conclusion, DSci requested UVCS, CELIAS, EIT, and CDS PI’s to report on 14 September 1994 at ESTEC on current technical and schedule status.

7 SOHO Intercalibration Committee report

SIC minutes — see annex 12 J. Kohl gave a report of the SOHO Intercalibration Committee (SIC) meeting which took place Monday afternoon (June 13, 14:00–18:00 hours).

8 SPWG report

A. Gabriel gave a report of the SOHO Science Planning Working Group (SPWG) meeting which took place on Tuesday, June 14 (see annex 13). In the morning session the progress in preparing Joint Operation Plans (JOPs) was reviewed and new JOPs as well as the procedure for JOP planning and control were discussed. In the afternoon session the timeline for cruise phase science operations was presented by P. Martens, followed by a presentation by PI team representatives of the plans for the first 30 days of nominal science operations. There is an action on the PI’s to comment and update the early operations timeline plan distributed by P. Martens, including the commissioning phase [first 30 days after HOI] and the first 30 days of nominal operations after commissioning (Action 12.1). Input to P. Martens before July 31.

9 Data Rights and Obligations

V. Domingo presented the SOHO general policy on data rights (cf. SOP Draft Issue 2.0, Section 3.5.3.2 on Science Data Access Rights) and Table 3.7 of SOP Draft Issue 2.0 (Data availablity throughout the SOHO mission as agreed by the SOHO PI’s). There is an action (12.3) on the PI’s to comment to this table and provide their team’s data rights and publication policy (to be included in the SOP as Annex E), due July 31, 1994. Annex 14 contains an illustration of the different views on data openess as well as SOI-MDI comments on data rights.

10 SOHO Data Archive

L. Sanchez presented the milestones of the SOHO Data Archive implementation at the EOF at GSFC (see annex 15). The US teams expressed some concerns that a facility as described might cause a degradation of the funding of the US PI investigations (see SOI-MDI comments on data archives, also attached to annex 15). J. Kohl suggested to call this “live” archive a “data library”. R. Howard and J. Kohl expressed some concerns regarding the access control management of the data in the SOHO archive.

11 European SOHO Data Archive

V. Domingo presented the implementation plan of a European SOHO Data Archive(s) (see annex 16). Following a brief discussion, the SWT endorsed the ESA plan to seek institutions interested in hosting the SOHO data in Europe (see Resolution in Section 1). The European archive(s) will be a copy of the EOF archive (data library) and adhere to the same access rules established for the EOF archive.

12 Early Operations

P. Martens gave a report on early operations timelining (see annex 17). Early operations now includes the cruise phase (≈ 120 days), the commissioning phase (S/C + P/L; 30 days), and the initial science operations (30 days). There is an action on the PI’s (12.1) to comment and update the early operations timeline plan. Input to P. Martens before July 31 (preferably in electronic form).

13 Jitter Timeline

P. Martens presented the course of action for jitter timelining workaround:

1. Collect jitter data on experiments and S/C (Project provided)

2. Input to EOF/ECS timelining S/W (ECS S/W available since May)

3. Check proposed JOPs for compatibility with jitter criteria – Need input of detailed experiment procedures)

4. If violations, try to modify JOPs – Estimate required flexibility of JOPs

5. Report back to SWT/splinter – Devise rules on priorities

14 Guest Investigators Programme

P. Martens presented the new draft 2.2 of the SOHO GI AO. The following changes have been applied since SWT-11:

• It has been emphasized that SOHO is a PI, not an observatory type of mission, and that accepted GI’s will be participating in PI teams, rather than simply use the data.

• Two modes of GI participation have been defined, in accordance with reactions received from the PI teams: one for the coronal, and one for the particle and helioseismology instruments. For a description of these modes I refer to the new Section 2.1 of the draft: “Philosophy of the Guest Investigator Programme”.

• The criterion of “30% of observing time” has been dropped completely from the text, except as a possible criterion for the Guest Investigator Selection Committee for determining a cut-off in accepting proposals.

• The term PI has been reserved for SOHO PI’s while for the PI of the Guest Investigation the term Leading Guest Investigator has been used. This ends the confusion commented upon by many of you.

• In the time table the date for the Issue of the Announcement of Opportunity is moved back to April 1, 1995.

• In the Foreword reference is made to the STSP Programme, and the basis for the GI Programme in the STSP AO.

• The description of the EOF has been rewritten to emphasize the scientific aspects, rather than the technical.

15 SOP

B. Fleck presented the draft version of SOP Issue 2.0 which was circulated to the PI’s and SOWG members in early June. SOP Issue 2.0 will be released end of August this year after incorporating the comments received from the PI’s and SOWG members. Other open issues which will be addressed in the new issue 2.0 are:

• EOF/EAF splitting

• European SOHO data archives

• New Annex E: PI team’s data rights and publication policies (cf. Action 12.3)

16 Workshops and Conferences

V. Domingo gave a report of the IACG workshop on “Solar Events and their Manifestation in Interplanetary Space and in Geospace” which took place in Sagamihara, Japan, 31 May to 2 June, 1994 (see annex 18). The 3rd SOHO Workshop on “Solar Dynamic Phenomena & Solar Wind Consequences” will take place in Estes Park, CO, USA, Sep. 26–29, followed by a 1-day SPWG meeting on Friday, Sep. 30. The 4th SOHO Workshop will be on Helioseismology and will take place in the Asilomar Conference Center in Pacific Grove, CA, April 2–6, 1995.

17 Science Working Team executive meeting

B. Fleck informed the PI’s about the agreement with SOLAR PHYSICS to have the June issue of 1995 for a SOHO special containing all the instrument descriptions. There is an action on the PI’s (12.4) to send camera ready (Kluwer format) instrument papers before 31 October 1994 to B. Fleck who will coordinate the editing with SOLAR PHYSICS. More detailed information will be circulated before end of July.

There are 4 action items and 1 resolution from SWT-12 (see Sect. 1).

The next meeting (SWT-13) will take place in the week of December 12 at ESTEC or Toulouse.

Annex 1: Agenda

8

Annex 2: Attendees SWT-12

Name  Institute/Experiment  Address

Bedini, P. Univ. of Maryland / CELIAS- Space Physics Group, College Park,

MTOF MD 20742 (bedini@umdspa.umd.edu)

Berner,C.  ESA / PLP  ESTEC (cberner@estec.esa.nl)
Bertaux, J.L.  Service d’Aeronomie / SWAN  BP 3, Verrieres-Le-Buisson, F-91371 (savera::bertaux)
Brueckner, G.E.  NRL / LASCO  Code 7666, Naval Research Lab,

Washington, DC 20375-5352, USA

Calabrese, M.  NASA / HQ  NASA Headquarters, Washington DC, USA
Charra, J.  IAS / GOLF  Univ. Paris XI, Bat 121, F-91405 Orsay Cedex
Defise, J.M.  EIT  CSL av. du Pre Aily, B-4031 Angleur

Delaboudiniere, J.P. IAS / EIT Bat 121, Univ. Paris XI, F-91405 Orsay Cedex

Domingo, V.  ESA / SSD  ESTEC (vdomingo@soho.estec.esa.nl)
Duc, R.  GOLF  Service d’Astrophysique, CEA Saclay, France
Felici, F.  ESA / PL  ESTEC (ffelici@estec.esa.nl)
Fleck, B.  ESA / SSD  ESTEC (bfleck@soho.estec.esa.nl)
Fröhlich, C.  PMOD-WRC / VIRGO  CH-7260 Davos-Dorf
Gabriel, A.  IAS / GOLF  Bat 121, Univ. Paris XI, F-91405 Orsay Cedex
Galvin, A.B.  Univ. of Maryland / CELIAS  Space Physics Group, Dept. of Physics,

College Park, MD 20742, USA

Gardelle, J.P.  MATRA  Matra Marconi Space, Toulouse, France
Harrison, R.A.  RAL / CDS  RAL, Didcot, Oxon, OX11 OQX, UK

(rah@uk.ac.rl.astro)

Hilchenbach, M.  MPE / CELIAS  MPE, D-85740 Garching
Hovestadt, D.  MPE / CELIAS  MPE, D-85740 Garching
Ipavich, F.M.  Univ. of Maryland / CELIAS  Space Physics Group, Dept. of Physics,

College Park, MD 20742, USA

Judge, D.L.  CELIAS / SEM  Univ. of Southern California
Kent, B.J.  RAL / CDS  RAL, Didcot, Oxon, OX11 OQX, UK
Kohl, J.  SAO / UVCS  SAO, 60 Garden ST, Cambridge, MA 02138, USA
de Koomen, J.H.  Ball Aerospace / UVCS  Nieuwe Bussummerweg 166, Huizen, 1272 CM, NL
Lemaire, P.  IAS / SUMER  Bat 121, Univ. Paris XI, F-91405 Orsay Cedex
Machi, D.  NASA / GSFC  GSFC, MD 20771, USA (dmachi@gsfcmail.nasa.gov)
Martens, P.  ESA / SSD  ESTEC (pmartens@soho.estec.esa.nl)
Michels, D.  NRL / LASCO  Code 7666, Naval Research Lab,

Washington DC 20375-5000, USA

Mueller-Mellin, R. Univ. of Kiel / CEPAC Otto-Hahn-Platz 1, D-24118 Kiel

Norman, K.  MSSL / UCL  Holmbury St. Mary, Dorking, Surrey RH56NT, UK
Poland, A.  NASA / GSFC  Code 6821 Greenbelt MD 20771, USA (pal::poland)

Sanchez-Duarte, L. ESA / SSD ESTEC (lsanchez@soho.estec.esa.nl)

Sawyer, E.  RAL / CDS  RAL, Didcot, Oxon, UK, OX11 OQX

(esu@uk.ac.rutherford.ibm-b)

Scherrer, P. Stanford University / SOI-MDI CSSA, Stanford, CA 94305-4055, USA

Sequeiros, J.  CEPAC  Univ. of Alcala, Madrid, Spain
Sizemore, K.  NASA / GSFC  GSFC, Code 407, Greenbelt MD 20771, USA
Smith, A.  MSSL / UCL  Holmbury St. Mary, Dorking, Surrey RH56NT, UK
Thomas, R.  ESTEC / PLQ  ESTEC (estcs1::rthomas)

Torsti, J.  University of Turku / ERNE  Space Res. Lab., SF-20520 Turku (torsti@sara.utu.fi)

Vandenbussche, F. ESA / PLS ESTEC (fvandeb@estec.esa.nl)

Wenzel, K.P.  ESA / SO  ESTEC (kwenzel@estec.esa.nl)
Wilhelm, K.  MPAE / SUMER  D-37189 Katlenburg-Lindau (nsp::linmpi::wilhelm)
Wolfson, C.J.  Lockheed / SOI-MDI  LPARL, 3251 Hanoverstr., Palo Alto, CA 94304, USA
Wurz, P.  Univ. of Bern / CELIAS  Sidlerstr.5, CH-3012 Bern (wurz@phim.unibe.ch)

Annex 3: Science Operations Report (V. Domingo)

11

Annex 4: EOF Layout (A. Poland)

14

Annex 5: SOHO CD-ROMs Distribution List (A. Poland)

15

Annex 6: Project Status (F. Felici)

20

Annex 7: SOHO Engineering Review (F. Vandenbussche)

27

Annex 8: Payload & Operations Development Status (C. Berner)

38

Annex 9: Cleanliness Review (R. Thomas)

74

Annex 10: Task Force Report (L. Culhane)

90

Annex 11: Status of Experiments

SOI-MDI (P. Scherrer)

GOLF (A. Gabriel)

116

VIRGO (C. Fröhlich)

122

LASCO (G. Brueckner)

128

CDS (R. Harrison)

140

SUMER (K. Wilhelm)

161

SWAN (J.-L. Bertaux)

177

CELIAS (D. Hovestadt)

184

COSTEP (R. Müller-Mellin)

195

ERNE (J. Torsti)

204

EIT (J.-P. Delaboudiniere)

214

UVCS (J. Kohl)

216

Annex 12: SIC Report (J. Kohl)

232

Annex 13: SPWG Report (A. Gabriel)

235

Annex 14: Data Rights and Obligations (V. Domingo)

240

Annex 15: SOHO Data Archive (L. Sanchez)

244

Annex 16: European SOHO Data Archive (V. Domingo)

254

Annex 17: Early Operations (P. Martens)

261

Annex 18: IACG Workshop Report (V. Domingo)

269